a machine can manufacture 1800 pencils in 30 minutes at this same rate how many minutes will it take to manufacture 3000 pencils
Answer by TimothyLamb(4379)

30/1800 = x/3000
x = 3000*30/1800
x = 50 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php